OpenStudy (anonymous):

A huricane is 300 miles offshore and moving at 50 mph. For what range of time will the hurricane be between 50 and 150 miles offshore. Write the lower and upper limits separated by a comma

OpenStudy (ivancsc1996):

\[t=\frac{ x }{ v }\rightarrow t _{150m}=\frac{ 150m }{ 50\frac{ m }{ h } }=3h, t _{50m}=\frac{ 250m }{ 50\frac{ m }{ h }}=5h\]This is a Physics question and should be asked on the physics section. The answer is 3h, 5h. Any questions?

OpenStudy (radar):

How much time will this hurricane traverse 100 miles (the distance of 150 to 50 is 100 miles) The storm is traveling at 50 mph, no mention is made on the width of the storm so the storm can be treated as a "front" moving for a distance of 100 mph. d=rt distance = rate (speed) times time (t) 100 = 50 t t = 100/50 = 2 hours. The range of time would be as follows making some assumptions, the hurricane is moving toward you so it will approach the 150 mile range in 3 hours, it will approach the 50 mile range in 5 hours (3,5) 5-3=2 hours. This question seems to be in the right place, as it is a mathematical question rather than a physics one.